# Structured Unrestricted-Rank Matrices for Parameter Efficient Fine-tuning

### v2

@@ 7 Conclusion
We introduce structured unrestricted-rank matrices (SURMs) as an alternative to low-rank matrices for the parameter efficient fine-tuning (PEFT) of large Transformer models. In this setting, structured matrices form the cornerstone of a comprehensive framework, offering a solid base for various PEFT methodologies, such as adapters and LoRA, with enhanced efficiency. SURMs improve the overall effectiveness of PEFT, contributing to its efficient integration into diverse models and domains. Based on extensive numerical experiments and theoretical insights, we conclude that the Circulant variant is our most performing variant (in terms of speed and accuracy).
@@ Appendix A Implementation Details
A.1 Finding the Smallest Number of Training Parameters for Kronecker Layers
@@
A.2 Efficient Matrix Vector Multiplication by Structured Matrices
@@
A.3 Integration of SURMs in Adapters
The adapter method [25] proposes injecting small bottleneck networks into Transformer layers, usually after feed-forward layers. The main equation of an adapter block is given by:
@@ A.3 Integration of SURMs in Adapters
SURMs can be used in place of low rank $\mathbf{A}$ and $\mathbf{B}$. The integration and design choices of various LDRs in this setting mimic that of LoRA. We will now provide a detailed explanation of the integration of circulant matrices within the adapter setting.
@@
where $f$ is an operator multiplying input matrix $\mathbf{X}$ with the circulant matrix obtained by multiplying two circulant matrices encoded by $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ (Appendix A.2). The vector $\mathbf{r}_{1}$ is initialized randomly while $\mathbf{r}_{2}$ and $\mathbf{b}$ are initialized as zero vectors. Note that we apply the non-linearity after we multiply $\mathbf{X}$ with both the circulant matrices. This may hurt the expressiveness of the network but improves computational complexity. Moreover, we only need to save one vector defining the first row of a circulant matrix and not both: $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$. This results in lower storage costs and faster deployment.
@@
Kronecker Product of Matrices. In this case, we rewrite Equation 6 as:
${}\mathbf{Y}=\mathbf{X}+\sigma(\mathbf{X}(\mathbf{B}\otimes\mathbf{A}))+\mathbf{b},$ (9)
where $\mathbf{B}\otimes\mathbf{A}$ is the Kronecker product. In this case, $\mathbf{B}$ is initialized randomly and $\mathbf{A}$ and $\mathbf{b}$ are initialized by zeros. In all experiments using SURM-adapters, $\sigma(\cdot)$ is the GeLU non-linearity.
A.4 Computing Approximations by LDR Matrices
@@
[…]
@@
A.5 Additional Details on Approximation Errors by LDR
@@
A.6 Invertible Toeplitz Matrices
@@ Appendix B Experiments
In this section, we describe our experimental setup and present additional analysis experiments to evaluate the functioning of SURM.
@@ B.1 Hyperparameters
Additional Details on the Pinwheel Experiment. We provide additional details on the pinwheel experiment. We tried out 2 settings : (a) simple neural network training for 2000 epochs, (b) the embedding (bottom) layer is frozen and the rest of the network is trained for 2000 epochs. This can be thought of fitting a feature extractor on top of a randomized projection. The setting (b) is presented in the main paper while setting (a) is presented in Appendix B.2. Next, we provide additional details for our text and vision experiments.
Figure 8: Left : Representation of the pinwheel dataset., Right : Fitting the same dataset with regular training.
@@ B.2 Additional Experiments
First, we provide a figure of the pinwheel dataset used to showcase the approximation qualities to LDRMs (see Fig 8 left). Next, we conduct additional experiments on the ImageNet-1k dataset [18]. The goal is to show how our methods can scale up to extremely large datasets. We observe that SURM achieves comparable performance to other PEFT methods and even achieving comparable performance to the full fine-tuning results.
@@
Ablation Experiments. Here, we show the effect of various design choices. Figure LABEL:fig:bias_ablates illustrates the impact of incorporating the bias term in our adapters. Bias term provides a boost across all tasks and the adapters, the boost being smaller on the Kronecker adapter. Without the bias terms, the sizes of the adapters are around .$\mathbf{.04}$M, providing an even lightweight but still capable method. Therefore, if there are concerns regarding storage and latency, opting for adapters without bias is a viable option. Moreover, we show the effect on only adapting $\mathbf{Q},\mathbf{V}$ instead of $\mathbf{Q},\mathbf{K},\mathbf{V}$ as shown in the main paper. Table 7 shows that on GLUE tasks, there is a minimal effect for not adapting the $\mathbf{K}$ matrix.
@@ B.4 Analysis of Weight Matrices in Fine-tuned Models
In this section, we analyze the weights of various fine-tuned models. Even though prior works have found the updates of the weight matrices to have low intrinsic dimension [1] (ID), the updates themselves are of high rank. This is confirmed by looking at the fine-tuned BERT models on various GLUE tasks as well as ViT models fine-tuned on CIFAR10, CIFAR100, and ImageNet. Moreover, we simulate a high-rank LoRA setting on GLUE where we freeze all the weights excepts except for $\mathbf{Q},\mathbf{K},\mathbf{V}$. In that scenario, we manage to replicate the full fine-tuning performance using fewer training epochs than that of LoRA. A quick analysis of the updates reveal that they have full rank.

### v3

@@ 7 Conclusion
We introduce structured unrestricted-rank matrices (SURMs) as an alternative to low-rank matrices for the parameter-efficient fine-tuning of large Transformer models. In this setting, structured matrices form the cornerstone of a comprehensive framework, offering a solid base for various parameter efficient fine-tuning methods, such as adapters and LoRA, with enhanced efficiency. SURMs improve the overall effectiveness of PEFT, contributing to its efficient integration into diverse models and domains. Based on extensive numerical experiments and theoretical insights, we conclude that the Circulant variant is our most performing variant (in terms of speed and accuracy).
@@ Appendix A Implementation Details
[…]
A.1 Skew-Circulant Matrices
In this section, we introduce another type of structured matrix that is characterized by a linear number of parameters.
Definition A.1 (Skew-Circulant).
A matrix $\mathbf{S}=(s_{jk})_{j,k=0}^{n-1}$ is said to be skew-circulant if $s_{jk}=s_{j-k}$ and $s_{-l}=-s_{n-l}$ for $11\leq l\leq n-$ 1.
This matrix can be represented visually as shown below:
$\scriptstyle{{\displaystyle\mathbf{S}={\begin{bmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-v_{n-1}}&\cdots&{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}-v_{1}}\\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&\cdots&{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}-v_{2}}\\ \vdots&\vdots&\vdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-v_{n-1}}\\ {\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}v_{n-1}}&\cdots&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}\\ \end{bmatrix}}}}$ (5)
This matrix is parameterized by a linear number of parameters and also enjoy sub-quadratic time complexity matrix-vector multiplications (see Appendix A.3).
A.2 Finding the Smallest Number of Training Parameters for Kronecker Layers
@@
A.3 Efficient Matrix Vector Multiplication by Structured Matrices
@@
Next, we describe how vector multiplication by skew-circulant matrices can be efficiently performed in $O(n\log n)$ time. If the skew-circulant matrix $\mathbf{S}$ is parameterized by the vector $\mathbf{v}$, then the multiplication is given by
$\mathbf{y}=\mathbf{S}(\mathbf{v})\mathbf{x}=\bar{\boldsymbol{\eta}}\circ\text{ifft}(\text{fft}(\boldsymbol{\eta}\circ\mathbf{v})\circ\text{fft}(\boldsymbol{\eta}\circ\mathbf{x}))$ (8)
where $\boldsymbol{\eta}=[1,\eta,\eta^{2},\cdots\eta^{n-1}]$, and $\eta=(-1)^{\frac{1}{n}}=\exp(i\pi/n)$, the root of negative unity.
@@
A.4 Increasing Number of Training Parameters
In this section, we explain an elegant way to increase the number of training parameters. We use the sum of product of circulant and skew-circulant matrices of the form
${}\mathbf{M}=\sum_{i=1}^{r}\mathbf{A}_{i}\mathbf{B}_{i}$ (9)
where $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$ is a circulant and a skew-circulant matrix respectively. Each factor of $\mathbf{M}$ has $2n$ parameters thus $\mathbf{M}$ has $2nr$ parameters.
The class of $n\times n$ matrices $\mathbf{M}$ which can be written via Equation 9 is rich and contains many important classes of matrices.
Theorem A.2 (Expressivity).
The set of matrices $\mathbf{M}$ which can be written as in Equation 9 contains:
  • All $n\times n$ Circulant and Skew-Circulant matrices for $r\geq 1$
  • All $n\times n$ Toeplitz and Inverses of Toeplitz matrices for $r\geq 2$.
  • All $n\times n$ matrices for $r=n$.
  • All linear combinations of the form $\sum_{j=1}^{p}\beta_{i}\mathbf{A}_{1}^{(j)}\cdots\mathbf{A}_{t}^{(j)}$ where $r\geq 2tp$, and $\mathbf{A}$ is either a Toeplitz or the inverse of a Toeplitz matrix.
This is Theorem 3.1 in [66]. Efficient multiplication by matrices of this form can be done in sub-quadratic time by simply combining the results from Sec A.3.
Moreover, we note that by choosing a slightly different parameterization of displacement operators, one can obtain low rank matrices and orthogonal polynomial transforms, including the Discrete Fourier and Cosine Transforms (see Proposition 2 in [68]).
Thus our framework encompasses many important classes of matrices including low rank matrices and thus generalizes LoRA.
A.5 Integration of SURMs in Adapters
In this section, we provide additional details about SURMintegration in Adapters.
For simplicity, we follow the Houlsby configuration [25], but our work is also readily applicable in Pfeiffer configuration as well [59]. Recall the definition of Adapter layers:
@@ A.5 Integration of SURMs in Adapters
SURMs can be used in place of low rank $\mathbf{A}$ and $\mathbf{B}$. The integration and design choices of various LDRs in this setting mimic that of LoRA.
@@
where $f$ is an operator multiplying input matrix $\mathbf{X}$ with the circulant matrix obtained by multiplying two circulant matrices encoded by $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ (Appendix A.3). The vector $\mathbf{r}_{1}$ is initialized randomly while $\mathbf{r}_{2}$ and $\mathbf{b}$ are initialized as zero vectors. Note that we apply the non-linearity after we multiply $\mathbf{X}$ with both the circulant matrices. This may hurt the expressiveness of the network but improves computational complexity. Moreover, we only need to save one vector defining the first row of a circulant matrix and not both: $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$. This results in lower storage costs and faster deployment. This design choice works well in practice as evidenced from the results on the GLUE benchmark (see Table 5).
@@
Finally, note that the Toeplitz variant is slower than the circulant variant, as it requires two applications of the fast matrix-vector operator, whereas the circulant variant requires only one.
A.6 Computing Approximations Using LDR Matrices
@@
The space of circulant matrices has dim $n$, so spanned by the orthogonal set $\{(1,\cdots,\cdots 0),(0,\cdots 1,\cdots 0),(0,\cdots,\cdots 1)\}$. Using the above formula, one can write down a simplified expression of the circulant matrix as $\hat{\mathbf{C}}:=(\hat{c_{1}},\cdots\hat{c_{n}})$ that approximates $\mathbf{D}$
@@
A.7 Additional Details on Approximation Errors by LDR
@@
A.8 Invertible Toeplitz Matrices
@@ Appendix B Experiments
In this section, we describe our experimental setup and present additional analysis experiments to evaluate the functioning of SURM. Our code is available at https://github.com/arijitthegame/structured-matrices-PEFT.
[…]
@@ B.1 Hyperparameters
Additional Details on the Pinwheel Experiment. First, we provide a figure of the pinwheel dataset used to showcase the approximation qualities to LDRMs (see Fig 9 left). We provide additional details on the pinwheel experiment. We tried out 2 settings : (a) simple neural network training for 2000 epochs, (b) the embedding (bottom) layer is frozen and the rest of the network is trained for 2000 epochs. This can be thought of fitting a feature extractor on top of a randomized projection. The setting (b) is presented in the main paper while setting (a) is presented in Appendix B.2. Next, we provide additional details for our text and vision experiments.
Refer to captionRefer to caption
Figure 9: Experiment on fitting the pinwheel dataset.  Left: Visualization of the pinwheel dataset.  Right: Results of fitting the pinwheel dataset using regular training, where all network parameters are trained. A network with a low-rank hidden layer matrix struggles to fit the data, while those with SURMmatrices achieve a successful fit.
@@
Large Scale Experiments.  In this experiment, we investigate if large ranks are needed for learning new tasks. To circumvent the pre-trained knowledge in the Transformers, following [29], we generate random 10K pairs of Universally Unique Identifiers (UUIDs), each pair comprising two UUIDs with 32 hexadecimal values. The task requires the LLM to generate the corresponding UUID based on the input UUID. We use LLaMA-2 7B as base model [69] for this experiment. For the LoRA setting, we apply rank $256$ matrices to only the linear layers in the attention layer. For MoRA, we use the same setting as in [29]. However, note that in [29], the authors apply adaptation to all linear layers.
For the SURM methods, we follow the same setting as applying the adaptation to only the attention layers. The number of factors of $\mathbf{W}(\mathbf{G},\mathbf{H})$ is chosen to be $4$ (i.e. same number of parameters as a rank $4$ matrix). This experiment also highlights that for effective transfer learning LoRA needs to be applied to all linear layers (which is well-known in the LLM community). In Fig. 8, we observe that LoRA and MoRA struggles to fit the data whereas our method converges.
@@ B.2 Additional Experiments
In this section, we provide additional experiments to showcase the efficacy of SURMs.
Experiments on Large Scale Data  Next, we conduct additional experiments on the ImageNet-1k dataset [18]. The goal is to show how our methods can scale up to extremely large datasets. We observe that SURM achieves comparable performance to other PEFT methods and even achieving comparable performance to the full fine-tuning results (see Table 6).
@@
We further evaluate the performance of SURM in a large-scale setting using the iNat2021 dataset [72], which contains over $\mathbf{2.7}$ million training images, $\mathbf{100K}$ validation images, and $\mathbf{500K}$ test images, spanning $\mathbf{10,000}$ species (classes). Fine-tuning a ViT model on this dataset achieves an accuracy of 69.98%, while SURM (circulant) achieves 69.01%. Notably, our method requires only $\mathbf{55K}$ parameters, compared to $86M$ for full fine-tuning, demonstrating its efficiency in parameter usage while maintaining comparable accuracy.
Figure 10: Figures showing the effect of using bias on various adapters on the GLUE tasks. The red dashed line is the full fine-tuning baseline which is almost $\mathbf{2000x}$ larger than our adapters.
Ablation Experiments. Here, we show the effect of various design choices. Figure 10 illustrates the impact of incorporating the bias term in our adapters. Bias term provides a boost across all tasks and the adapters, the boost being smaller on the Kronecker adapter. Without the bias terms, the sizes of the adapters are around .$\mathbf{.04}$M, providing an even lightweight but still capable method. Therefore, if there are concerns regarding storage and latency, opting for adapters without bias is a viable option. Moreover, we show the effect on only adapting $\mathbf{Q},\mathbf{V}$ instead of $\mathbf{Q},\mathbf{K},\mathbf{V}$ as shown in the main paper. Table 7 shows that on GLUE tasks, there is a minimal effect for not adapting the $\mathbf{K}$ matrix.
@@ B.4 Analysis of Weight Matrices in Fine-tuned Models
In this section, we analyze the weights of various fine-tuned models. Even though prior works have found the updates of the weight matrices to have low intrinsic dimension [1] (ID), the updates themselves are of high rank. This is confirmed by looking at the fine-tuned BERT models on various GLUE tasks as well as ViT models fine-tuned on CIFAR10, CIFAR100, and ImageNet. Moreover, we simulate a high-rank LoRA setting on GLUE where we freeze all the weights excepts except for $\mathbf{Q},\mathbf{K},\mathbf{V}$. In that scenario, we manage to replicate the full fine-tuning performance using fewer training epochs than that of LoRA. A quick analysis of the updates reveals that they have full rank.
@@ Appendix C Broader Impact & Limitations
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide extensive empirical evidence in Section 6 as well additional motivating experiments in Figure 4
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work is detailed in Appendix C.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not contain any new theoretical results but builds on the main theorem in [66].
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the hyperparameters to replicate the experiments in Appendix B.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide access to the code on building the adapter layers using SURM and their integration into pretrained transformers.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the hyperparameters to replicate the experiments in Appendix B.1 and B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Given the sheer volume of experiments along with the baselines that we had to replicate, it was too expensive for us to run these experiments for multiple seeds. However, we report error bars for the smaller toy experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide this information in Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All authors have reviewed NeurIPS Code of Ethics and the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of our work is detailed in Appendix C.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We are not releasing any new data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the authors who created the original data and algorithms that are used in this work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the main code detailing the creation of the SURM layers leveraging fast matrix-vector multiplication and their integration in pretrained Transformers.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not use any crowd sourcing and human subjects. The multi-organ segmentation downloaded from synapse and we adhere to the rules of the usage of this data.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not IRB approval for the use of the multi-organ segmentation dataset. This can be freely downloaded from the internet once one signs up on synpase. We strictly adhere to the rules of the usage of data.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.